OpenStudy (anonymous):

Find the derivative of the function. y = 8^( 6-x^2 )

OpenStudy (anonymous):

I know to use the chain rule but I can't figure out how to apply it. Please help?

OpenStudy (abb0t):

I would use logrithmic differentiation for this. then product rule:\[\ln(a)^b = b \times \ln(a)\]

OpenStudy (anonymous):

 is this correct so far?